OpenStudy (nicole143):

Binomial distribution with the expression (q + p)^n where n = 5 and p = 0.3 Can you please explain how to do this?

OpenStudy (nicole143):

@kirbykirby @nincompoop Help please?

OpenStudy (kirbykirby):

Um are you trying to expand this using the binomial theorem? I'm not sure what you need since you mention the binomial distribution.

OpenStudy (nicole143):

All it say is to use the binomial expression to calculate the binomial distribution and I don't know how to do so..

OpenStudy (kirbykirby):

Do you have the full question? Do they just want you to find the probability mass function (pmf)? The binomial distribution is: \[ f_X(x)={n\choose x}p^xq^{n-x}\] where \(q=1-p\) If you're given p and n, then you know q... and plug in those values... then the probability of an event X=x is given by the pmf above.

OpenStudy (nicole143):

The full question is: Use the binomial expression (q + p)^n to calculate a binomial distribution with n = 5 and p = .3

OpenStudy (kirbykirby):

Ok then I think that's what you need then. So you can calculate each probability, then you plug in values for \(x=0,1,\ldots,5\). (Notice that \(q=1-p=1-0.3 =0.7\)) So for example: at \(x=0:\) \[f_X(0)={5\choose 0}0.3^{5}0.7^{5-0}\] at \(x=1:\) \[f_X(1)={5\choose 1}0.3^{5}0.7^{5-1}\] .... until \(x=5\) Just need to calculate those values on your calculator

OpenStudy (kirbykirby):

The confusing part of the question is that I don't know why they give you a binomial (p+q)^n. I don't see why that is needed to compute a binomial \(\textit{distribution}\), especially since that binomial expression just evaluates to 1...

OpenStudy (nicole143):

x = 1 1,221.205 ??

OpenStudy (kirbykirby):

No, each value will be between 0 and 1.

OpenStudy (nicole143):

Oh. Okay, will you look at this? https://answers.yahoo.com/question/index?qid=20140325093333AAzGSpG Is this what I have to do?

OpenStudy (kirbykirby):

The 2nd answer is basically what I described, but they are using the notation 5C1 instead of \(\large {5 \choose 1}\).

OpenStudy (kirbykirby):

The 1st answer they just did a binomial expansion, but this is not finding a binomial distribution.

OpenStudy (nicole143):

Okay, can yo take me through it step by step?

OpenStudy (kirbykirby):

Um.. I don't know what to say other than you just need to plug in the given values, \(n\) and \(p\) into the binomial distribution distribution: \[f_X(x)={n\choose x}p^xq^{n-x}\] And of course, \(q=1-p\), since \(p\) is the probability of success, and we know that the sum of all probability events sum to 1. You have either a success, or a failure (whose probability is represented by \(q\)), so \(p+q=1\) Maybe you're unsure how to calculate the binomial coefficient? That is the \(\large {n \choose x}\) term?

OpenStudy (nicole143):

I am unsure of how to calculate that term..

OpenStudy (nicole143):

Can you also explain the 5-0 that is the power of 7 ?

OpenStudy (nicole143):

Do i just do 5, 4, 3, 2, 1 for x = 1 and x = 2 all the way to 5?

OpenStudy (kirbykirby):

\[ {n \choose x} =\frac{n!}{x!(n-x)!}\] So say you have \(n=8, x=3\): \[ {8\choose 3}=\frac{8!}{3!(8-3)!}=\frac{8!}{3!\cdot 5!}\] Here, \(8!=8\times 7 \times 6 \times 5 \times 4 \times 3\times 2 \times 1\) Similarly, \(5!=5 \times 4 \times 3\times 2 \times 1\) \(3!=3 \times 2\times 1\) So: \[\frac{8!}{3!\cdot 5!}=\frac{8\times 7 \times 6 \times \cancel{5 \times 4 \times 3\times 2 \times 1}}{(3 \times 2\times 1)\cdot\cancel{(5 \times 4 \times 3\times 2 \times 1)}}=\frac{8\times 7 \times 6}{3\times 2}=56\]

OpenStudy (nicole143):

Oh, okay. What about my other two question?

OpenStudy (nicole143):

x = 0 for the first one?

OpenStudy (kirbykirby):

\[f_X(0)={5\choose 0}0.3^{5}0.7^{5-0}\]... Here I just substituted the \(x\) with 0. Be careful, here you have the decimal numbers 0.3 and 0.7 raised to a power. To make it more clear: \[f_X(0)={5\choose 0}(0.3)^{0}(0.7)^{5-0}\] I should point out at this stage that ... I made a typo when I inputted "5" in the exponent of 0.3 Sorry about that =( There is clearly an \(x\) there in the formula. I just left the 5-0 there, rather than just 5, just to show you what you need to do in the formula. Now for x=1: just do: \[f_X(1)={5\choose 1}(0.3)^{1}(0.7)^{5-1}=5(0.3)(0.7)^4\] Now for x=2: just do: \[f_X(2)={5\choose 2}(0.3)^{2}(0.7)^{5-2}=10(0.3)^2(0.7)^3\]

OpenStudy (nicole143):

Okay, is this right?

OpenStudy (kirbykirby):

yes looks good, although in the first line you wrote \(\large {5 \choose 0}\) even though you still correctly put 1 instead of 0 in subsequent steps ;)

OpenStudy (nicole143):

Okay. Ha thanks, dang so now I have to do all that 4 more times?

OpenStudy (kirbykirby):

Yeah. but also include the case for x=0.

OpenStudy (nicole143):

Alright well thank you very much! Just one more thing, at the very end after I do all the rest do I need to write 1 through 5 in just answer form? Could you hope me with one more?

OpenStudy (nicole143):

*help

OpenStudy (kirbykirby):

I suppose you could? I'm not sure how detailed your instructor needs your answers to be :S

OpenStudy (kirbykirby):

Sure

OpenStudy (nicole143):

Ha okay. I'll open it in another post so you get both medals :P

OpenStudy (kirbykirby):

Hehe ok